Calculate the bulk modulus of air from the following data about a sound wave of wavelength 35 cm travelling in air. The pressure at a point varies between (1.0 × 106 ± 14) Pa and the particles of the air vibrate in simple harmonic motion of amplitude 5.5 × 10-6 m.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. We know that, Bulk modulus B = ∆p/(∆V/V) = p base λ/2πs base 0 Where P base 0 = pressure amplitude ⇒ P base 0 = 1.0 × 10^5 S base 0 = displacement amplitude ⇒ S base 0 = 5.5 × 10^–6 m